Weekly LIPAc highlights (27/Sepember-1/October)

As you may remember the LIPAc beam operation ended at the end of July.

Starting from August and until end of September there are major maintenance interventions in LIPAc, including regulatory inspections (on electrical system, cooling systems, HVAC, crane in the vault etc.) compulsory by law.


The new Emittance Meter Unit delivered by F4E has been installed in the vault, cleaned and the compatibility with the ancillaries and mechanical connections confirmed.

The main maintenance on the injector has just started and will continue until early October.

Injector conditioning is planned to resume by mid of October.

Radio Frequency (RF) System and RFQ cavity

The full maintenance of HVPS and RF-RFQ modules was performed with the support of NAT subcontractor. The RF team is currently working on upgrades of the RFQ RF modules to improve their functionality and stability in operation: upgraded PCB cards for the anode boxes have been validated and are being installed; PSYS VSWR cards to fit the new predrivers are under calibration; AC filters have been installed inside the PLC for noise reduction purposes.

By mid-October the RFQ RF modules will resume operation and target end of October to start injection in the RFQ. RF team will then need one week to perform three activities that needs RF injection in the cavity and then the RF will be ready for beam operation.

Beam Transport

Maintenance, checks and calibration of MEBT and MEL magnets is almost completed. HEBT magnets calibration is planned next week.

A new radiation resistance RGA and a new rupture disk will be installed this week in HEBT vacuum sector IV. The calibration and cable routing for the RGA are completed.


A part the standard maintenance, several interventions were performed in the cooling skids (active carbons replacement and UPS replacement in the RF skid, valves inspections and recovery of functionality). Some pending interventions on the skids requires specific skills and a dedicated contract is being prepared by QST targeting the execution at the end of the year.

Also the newly commissioned Beam Dump skid underwent a thorough maintenance in September.


Remote sessions with CIEMAT support have resumed to complete the commissioning of the HEBT and Beam Dump instrumentation (Faraday Cup, Fluorescence Profile Monitors, SEM grids motion control), debug HEBT PLC and complete the commissioning of the HEBT vacuum machine state. The pending hardware and cabling installation work was mostly executed onsite in August and remote sessions have been resumed in September

A VME control issues has to be solved to allow completing all the checks on Current Transformers and SEM grids data acquisition. CEA is supporting remotely.

A BPM in the MEBT is still not fully operational: the plan to remove the upper half of the quadrupole (MEBT BPM are embedded in the magnets) to access the BPM has been postponed to avoid the risk of the intervention.


In September a large Helium buffer tank was installed using large cranes (see pictures below).

André Duarte was on-site for a month in July-August (first mission after a long time) to work on the integration in EPICS of the cryoplant (this included installation of a new PLC in the cubicles, connection to the local net and debugging).

Control system.

The Personal Protection system (PPS) annual maintenance and tests were carried out several weeks in September and are completed.

In the current Integrated LIPAc schedule, it is planned to resume beam operation early November.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.